OpenStudy (anonymous):

A line has a slope of zero and passes through the point (-2, 3). Which of the following points must also lie on the line?

OpenStudy (anonymous):

Where are the following points?

OpenStudy (texaschic101):

if it has a slope of zero, then it is a horizontal line. One point on the line would be (anything for x, and 3 for y).

OpenStudy (texaschic101):

a horizontal line is represented by y = a number....that number is 3...it will always be 3. Your x can be any number, buy your y value, on this particular problem, is always going to be 3

OpenStudy (anonymous):

that's all they give me and I have 3 possible answers

OpenStudy (anonymous):

1- (-2, 5) 2- 5, 3) 3- (2, -3)

OpenStudy (texaschic101):

okay...look at your answer choices. your points are set up like (x,y). Now which one has a y value of 3 ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

answer 2

OpenStudy (texaschic101):

BINGO...you got it natasha...answer 2....good job :)

OpenStudy (anonymous):

thank you can you help with one more please

OpenStudy (texaschic101):

Here is something that might help you... HOY -- stands for : Horizontal line, 0 slope, y = a number VUX -- stands for : Vertical line, undefined slope, x = a number oh...I can try to help

OpenStudy (anonymous):

this one its true or false question The set of points C(1, -1), D(3, 4), E(5, 8) are collinear.

OpenStudy (texaschic101):

well....colinear means they are on the same line... lets make a table.. x 1 3 5 ----: --- : --- : ---- y -1 4 8 so when x is 1, y = -1....when x is 3, y = 4...so far, the difference in the x's is 2 and the difference in the y's is 5....so for it to be on the same line, the difference have to be the same. 3 + 2 = 5 (correct) 4 + 5 = 9 (incorrect)...for it to be on the same line, it would have to have been 5/9 FALSE

OpenStudy (anonymous):

thank you so much that was a lot of help

OpenStudy (texaschic101):

sorry it took so long

OpenStudy (anonymous):

that's ok